results of surveying 200 high school seniors to find out how many hours they worked per week at a job. pie charts: 42 (15 hours or less) 82(more than 15 hours) 76 (Didn't work0
a. find the ratio of students who work more than 15 hours a week to total students
answers: 82/200=41/100
b. find the ratio of students who don't have a job to students who work more that 15 hours a week. answers: 76/200= 38=41
c. find the ratio of students with jobs to total students.
d. find the ratio of students with jobs to students without jobs.
I do not know how to solve problems c and d please help me.

Answer by drk(1908)

200 total students
students with jobs = 42 + 82 = 124
students without jobs = 76
----
c. find the ratio of students with jobs to total students.
answer
124/200 = 62/100 = 31/50
-----
d. find the ratio of students with jobs to students without jobs.
124/76 = 31/19